Find one ordered pair (x,y) of real numbers such that x + y = 8 and x^3 + y^3 = 200 - 4x^2 - 4y^2.
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



To help me get started on this, I first graphed the system of equations using desmos.com.

That graph showed that there are no real number solutions, so it would be a waste of time trying to solve the system algebraically.

ANSWER: no real solutions


Answer by ikleyn(52781)

.
Find one ordered pair (x,y) of real numbers such that x + y = 8 and x^3 + y^3 = 200 - 4x^2 - 4y^2.
~~~~~~~~~~~~~~~~~~~~~~~

Left side  x^3 + y^3  can be composed  as  (x+y)*(x^2 - xy + y^2).

The factor  (x+y)  can be replaced by 8,  based on the first equation.


After that, second equation can be written in this form

    8(x^2 - xy + y^2) = 200 - 4(x^2 + y^2),

or

    12(x^2 + y^2) - 8xy = 200,

    12(x+y)^2 - 24xy - 8xy = 200,

    12(x+y)^2 - 32xy = 200.


Again, we can replace here  (x+y)  by 8, and we get then

    12*8^2 - 32xy = 200,

    32xy = 12*64 - 200 = 568,

      xy = 568/32 = 17.75.


But under the condition x+y = 8,  well known minimax property says 
that the product xy may have the maximum value of  4*4 = 16  at  x = y = 8/2 = 4.


In other words, under given conditions, the original system has no real solutions.


So, there is no any pair of real numbers (x,y) satisfying the imposed conditions.

Solved completely in Math terms.